OpenStudy (samara8954):

For which operations is the set {–1, 1} closed? Choose all answers that are correct. A. addition B. subtraction C. multiplication D. division

OpenStudy (samara8954):

@ganeshie8

OpenStudy (anonymous):

multiplication only

OpenStudy (samara8954):

are you sure?

OpenStudy (mimi_x3):

actually no

OpenStudy (anonymous):

let me check one more time

OpenStudy (mimi_x3):

Ok well addition its obvious its not closed since 1+1=2 Subtraction its not closed since -1-1=-2 Division is not closed since 1/0=undefined So multiplication is the only one that is closed

OpenStudy (anonymous):

ok so you clarified mine so just multiplication

OpenStudy (mimi_x3):

Hmmm is this set include all real numbers or just integers?

OpenStudy (mimi_x3):

Anyways ya its just multiplication

OpenStudy (samara8954):

Which equations show that the set of whole numbers is not closed under subtraction? Choose all answers that are correct. A. 1 – (–2) = 3 B. 1 – 2 = –1 C. 2 – 0 = 2 D. 2 – 4 = –2

OpenStudy (freckles):

but @Mimi_x3 we don't have to worry about 0 since it is not in {-1,1}

OpenStudy (freckles):

so it is both division and multiplication

OpenStudy (anonymous):

well all we have is 1 and -1 so it would be just multiplication

OpenStudy (mimi_x3):

Welll freckles if its an interval then it includes 0 Thats my question ... is this an interval or is this just a set with 2 values including 1 and -1?

OpenStudy (freckles):

{-1,1} isn't an interval

OpenStudy (mimi_x3):

ohhhh i seee {-1,1} it just intervals ... not wearing my glasses

OpenStudy (mimi_x3):

K thanks :)

OpenStudy (anonymous):

so its both multiplication and division?

OpenStudy (mimi_x3):

Yes

OpenStudy (anonymous):

OK that would work i guess

OpenStudy (freckles):

yes because (-1)(1)=-1 and 1(1)=1 and (-1)(-1)=1 and 1(-1)=-1 all of these outputs are in {-1,1} just like we take any two numbers from the set and do all the possible divisions all the outputs will also be in {-1,1} like so: -1/1=-1 and 1/1=1 and -1/-1=1 and 1/-1=-1

OpenStudy (mimi_x3):

since 1/1=1 1/-1=-1

OpenStudy (anonymous):

oh i understand know,thanks!

OpenStudy (freckles):

if it was {-1,0,1} you guys would have nailed it

OpenStudy (anonymous):

oh yeah that would make scenes why i thought it was that.i had that question on a test 2 days ago.